# TabDPT: Scaling Tabular Foundation Models on Real Data

### v2

@@ 4.4 Ablation Study
In this section, we discuss the ablation of key components in our training and inference strategies, with results visualized in Figure 4.
Training Ablation   First, we assess the importance of SSL during training. To ablate SSL, we only use the original target for each table during training and observe a large loss in performance, as shown under “Supervised Target (Tr)” in Figure 4(a). The impact on training is further illustrated by the “Real-No SSL” curve in Figure 4(b). Training without SSL starts to overfit after around 50 epochs whereas with SSL model continues to improve even after 500 epochs (“Real - SSL” curve). These results demonstrate the critical role of SSL when training TFMs with real data where only one target is typically available per dataset.
Second, we assess the benefit of using retrieval during training to form the context versus random subsampling, the results are shown under “No Retrieval (Tr)” in Figure 4(a). We see that removing training retrieval leads to a consistent drop in both classification and regression test accuracy. Although smaller in magnitude than drop from removing SSL, these results confirm that aligning training and inference procedures is beneficial.
Finally, we benchmark the impact of training on real (“Real - SSL” curve) vs synthetic data (“Synthetic” curve) in Figure 4(b). We use the TabPFN [28] synthetic data generator for this experiment. We see that using real data with SSL consistently outperforms training on synthetic data across all epochs, achieving lower test loss under the same compute budget. This further highlights the effectiveness of real data when paired with SSL on our TFM architecture.
Inference Ablation   Similarly to [54], we find that using subsampling instead of retrieval during inference significantly decreases performance as indicated by “No Retrieval (Inf)” in Figure 4(a). Using a smaller context of 256 size also decreases performance as expected, although it does not decrease nearly as much as the other important components discussed above.
Figure 4: (a) Ablation of key components in training (Tr) and inference (Inf). A higher blue bar and a higher green bar indicate greater reduction in AUC and $R^{2}$ respectively. (b) Test loss curves on CC18 when training with and without SSL on real data as well as synthetic data only.
@@ 4.5 Scaling Laws
In Figure 1, we illustrate the scaling behaviour of our models along with the power-law fit. Since we train on equal proportions of classification and regression tasks, the loss on the $y$-axis represents the mean of the cross-entropy loss for classification and $1-\rho$ for regression, where $\rho$ is the correlation between prediction and target, equivalent to the MSE for normalized vectors. Visualization is done on a log scale for the excess loss ^$\hat{\ell}(P,D)-E$ instead of the raw loss, debiasing the estimate by $E$. We observe consistent improvements when both data and model size increase indicating that information contained in real world tabular datasets can be effectively mined with SSL to pre-train robust TFMs.
@@ 4.7 Inference Speed
In this section, we analyze the inference runtime of TabDPT against baselines on new datasets. For TFMs we measure the cost of computing the context and making forward passes through the models. For deep learning and tree-based methods that are dataset specific we measure the total time to train the model, including hyperparameter search, and to run inference. We repeat each experiment across dataset folds and measure the average times to process 1,000 rows (computed overall on $\approx 2M$ rows). For TabDPT we report runtimes with different context sizes from 128 to 2048 and the corresponding impact on accuracy. Results are shown in Figure 3(b), we see that even our largest model with context size 2048 is at least one order of magnitude faster than dataset-specific tree and deep learning baselines. TabDPT runtime is also comparable to the leading TFM TabPFN v2 while achieving higher accuracy. The cost of pre-training TFMs is not included in this comparison. However, analogous to LLMs, we stipulate that it is a one-time cost that is offset when model is applied across many datasets and use cases.
@@ 5 Conclusion and Future Work
We introduce TabDPT, an open tabular foundation model with a demonstrated ability to generalize on a range of unseen tabular datasets without additional tuning or adaptation. Our training approach provides an effective way to leverage real data with SSL to build robust TFMs. Models pre-trained with our procedure exhibit scaling laws with consistent improvements from both data and model size, analogous to foundation models in other domains. Given the practical ease of use and broad applicability of foundation models, we believe that these contributions will advance their adoption as an alternative to individually trained models in tabular domains. Future work involves incorporating dataset metadata such as column names and descriptions into the model to enrich representations, as well as extending TabDPT to the timeseries domain.
While TabDPT demonstrates strong performance, opportunities remain to further enhance TFMs and address current limitations. (i) Preliminary experiments using feature name embeddings led to overfitting. Expanding training data with more diverse tables containing free-form text may mitigate this and improve textual integration. (ii) TabDPT is designed for rectangular datasets with i.i.d. rows and does not explicitly model temporal dependencies, distribution shifts, or hierarchical structures. Methodological improvements could help overcome these constraints; e.g., ideas similar to Hoo et al. [31] can help tackle temporal dependencies. (iii) TabDPT could be enhanced with techniques from works such as Ma et al. [40] and van Breugel et al. [56] to complement its discriminative capabilities with generative modeling and density estimation.
@@ Appendix B Training Datasets
Figure B.1 provides a summary of the sizes and domains of the training datasets, and Table B.1 provides a full list of the datasets. Note that 93 datasets have classification targets, 29 datasets have regression targets, and 1 does not have a default target defined. However, we generate both classification and regression targets for each dataset by applying the SSL procedure described in Section 3.
@@
$1$ KDDCup09_appetency 50000 230 12M 61.9 Class. Human behaviour
@@
1479 hill-valley 1212 100 122K 0.0 Class. Deterministic and simulated
@@ C.1 Architecture Details
The model architecture (see Figure 2(b)) is comprised of input embedding functions, multiple transformer encoder layers, and task-specific output heads. The key architectural parameters are summarized in Tables C.1 and C.2.
@@
Preprocessing   We deliberately do minimal pre-processing of the data to ensure that our approach has wide applicability. All columns containing non-numerical values are mapped to integers using scikit-learn’s [45] LabelEncoder function. The table is then standardized to 00 mean and unit variance, and outliers beyond $10$ are clipped. After retrieval, we obtain a local context $X_{\text{ctx}}$ and their labels $y_{\text{ctx}}$. $X_{\text{ctx}}$ is standardized before the forward pass to avoid distribution shifts, and $y_{\text{ctx}}$ is also standardized for the same reason if it is a regression target.
@@ Appendix E Elo and Glicko2 Ratings
In Figures 1(a) and 1(b), we report the Elo and Glicko2 scores respectively. The results are consistent between the two plots, with TabDPT performing best on both metrics, followed by the leading TFM baseline TabPFN v2.
@@ F.1 Additional Results by Dataset Statistics
F.2 Results for IQM Estimator on CC18 and CTR23
Results for the raw scores using the InterQuantile Mean estimator using bootstrapping [1] are shown in Table F.1. The relative ordering and takeaways are the same as 1.
Algorithm CC18 CTR23
AUC Accuracy Correlation $R^{2}$
TabDPT 0.976 ± [0.974, 0.978] 0.928 ± [0.926, 0.931] 0.920 ± [0.918, 0.922] 0.847 ± [0.843, 0.851]
TabPFN v2 0.972 ± [0.970, 0.974] 0.917 ± [0.915, 0.919] 0.917 ± [0.911, 0.921] 0.841 ± [0.831, 0.848]
TabPFN (kNN) 0.959 ± [0.956, 0.962] 0.884 ± [0.881, 0.887] N/A N/A
TabPFN 0.939 ± [0.935, 0.943] 0.852 ± [0.849, 0.856] N/A N/A
TabR 0.967 ± [0.965, 0.969] 0.923 ± [0.920, 0.926] 0.909 ± [0.905, 0.912] 0.825 ± [0.817, 0.831]
MLP-PLR 0.967 ± [0.965, 0.968] 0.914 ± [0.911, 0.917] 0.907 ± [0.904, 0.910] 0.827 ± [0.822, 0.832]
MLP 0.915 ± [0.909, 0.920] 0.865 ± [0.860, 0.870] nan ± [nan, nan] nan ± [nan, nan]
XGBoost 0.965 ± [0.963, 0.967] 0.910 ± [0.906, 0.913] 0.904 ± [0.900, 0.907] 0.820 ± [0.814, 0.825]
LightGBM 0.964 ± [0.962, 0.967] 0.906 ± [0.902, 0.909] 0.900 ± [0.896, 0.904] 0.809 ± [0.803, 0.815]
CatBoost 0.964 ± [0.962, 0.967] 0.908 ± [0.905, 0.910] 0.897 ± [0.890, 0.903] 0.802 ± [0.794, 0.810]
kNN N/A N/A N/A N/A
Table F.1: Results on CC18 and CTR23. We report four metrics and their $95\%$ confidence intervals. The best algorithm for each metric is bolded.

### v3

@@ 4.4 Ablation Study
In this section, we discuss the ablation of key components in our training and inference strategies, with results visualized in Figure˜4.
Training Ablation   First, we assess the importance of SSL during training. To ablate SSL, we only use the original target for each table during training and observe a large loss in performance, as shown under “Supervised Target (Tr)” in Figure˜4(a). The impact on training is further illustrated by the “Real-No SSL” curve in Figure˜4(b). Training without SSL starts to overfit after around 50 epochs whereas with SSL model continues to improve even after 500 epochs (“Real - SSL” curve). These results demonstrate the critical role of SSL when training TFMs with real data where only one target is typically available per dataset.
Second, we assess the benefit of using retrieval during training to form the context versus random subsampling, the results are shown under “No Retrieval (Tr)” in Figure˜4(a). We see that removing training retrieval leads to a consistent drop in both classification and regression test accuracy. Although smaller in magnitude than drop from removing SSL, these results confirm that aligning training and inference procedures is beneficial.
Finally, we benchmark the impact of training on real (“Real - SSL” curve) vs. synthetic data (“Synthetic” curve) in Figure˜4(b). We use the TabPFN [31] synthetic data generator for this experiment. We see that using real data with SSL consistently outperforms training on synthetic data across all epochs, achieving lower test loss under the same compute budget. This further highlights the effectiveness of real data when paired with SSL on our TFM architecture.
Inference Ablation   Similarly to [58], we find that using subsampling instead of retrieval during inference significantly decreases performance as indicated by “No Retrieval (Inf)” in Figure˜4(a). Using a smaller context of 256 size also decreases performance as expected, although it does not decrease nearly as much as the other important components discussed above.
Figure 4: (a) Ablation of key components in training (Tr) and inference (Inf). A higher blue bar and a higher green bar indicate greater reduction in AUC and $R^{2}$ respectively. (b) Test loss curves on CC18 when training with and without SSL on real data as well as synthetic data only. We note that “Synthetic” and “Real - SSL” both saturate at 500 epochs, which we confirmed by training for 1024 epochs.
@@ 4.5 Scaling Laws
In Figure˜1, we illustrate the scaling behaviour of our models along with the power-law fit. Since we train on equal proportions of classification and regression tasks, the loss on the $y$-axis represents the mean of the cross-entropy loss for classification and $1-\rho$ for regression, where $\rho$ is the correlation between prediction and target, equivalent to the MSE for normalized vectors. Visualization is done on a log scale for the excess loss $\hat{\ell}(P,D)-E$ instead of the raw loss, debiasing the estimate by $E$. We observe consistent improvements when both data and model size increase indicating that information contained in real world tabular datasets can be effectively mined with SSL to pre-train robust TFMs.
@@ 4.6 Few-Shot Learning
Concurrently, Li et al. [40] found that TFMs performed strongly in the semi-supervised setting using a method distinct to ours; combining their approach with TabDPT is a promising avenue for future work.
@@ 4.7 Inference Speed
In this section, we analyze the inference runtime of TabDPT against baselines on new datasets. For TFMs we measure the cost of computing the context and making forward passes through the models. For deep learning and tree-based methods that are dataset specific we measure the total time to train the model, including hyperparameter search, and to run inference. We repeat each experiment across dataset folds and measure the average times to process 1,000 rows (computed overall on $\approx 2M$ rows). For TabDPT we report runtimes with different context sizes from 128 to 2048 and the corresponding impact on accuracy. Results are shown in Figure˜3(b), we see that even our largest model with context size 2048 is at least one order of magnitude faster than dataset-specific tree and deep learning baselines. TabDPT runtime is also comparable to TabPFN v2 while achieving higher accuracy. The cost of pre-training TFMs is not included in this comparison. However, analogous to LLMs, we stipulate that it is a one-time cost that is offset when model is applied across many datasets and use cases.
On the other hand, when training cost is not part of the budget (e.g., fixed models on fixed datasets), classical tree-based pipelines can deliver substantially faster inference (e.g., $\sim$0.02–0.10s per 1,000 rows), often about an order of magnitude faster than TabDPT for inference alone.
@@ 5 Conclusion and Future Work
We introduce TabDPT, an open tabular foundation model with a demonstrated ability to generalize on a range of unseen tabular datasets without additional tuning or adaptation. Our training approach provides an effective way to leverage real data with SSL to build robust TFMs. Models pre-trained with our procedure exhibit scaling laws with consistent improvements from both data and model size, analogous to foundation models in other domains. Given the practical ease of use and broad applicability of foundation models, we believe that these contributions will advance their adoption as an alternative to individually trained models in tabular domains.
While TabDPT demonstrates strong performance, opportunities remain to further enhance TFMs – including TabDPT – and address current limitations. (i) Preliminary experiments using feature name embeddings led to overfitting. Expanding training data with more diverse tables containing free-form text may mitigate this and improve textual integration. (ii) TabDPT is designed for rectangular datasets with i.i.d. rows and does not explicitly model temporal dependencies, distribution shifts, or hierarchical structures. Methodological improvements could help overcome these constraints; e.g., ideas similar to Hoo et al. [34] can help tackle temporal dependencies. (iii) TabDPT could be enhanced with techniques from works such as Ma et al. [44] and van Breugel et al. [60] to complement its discriminative capabilities with generative modeling and density estimation.
@@ Appendix B Training Datasets
Figure˜B.1 provides a summary of the sizes and domains of the training datasets, and Table˜B.1 provides a full list of the datasets. Note that 93 datasets have classification targets, 29 datasets have regression targets, and 1 does not have a default target defined. However, we generate both classification and regression targets for each dataset by applying the SSL procedure described in Section˜3.
@@
$11$ KDDCup09_appetency 50000 230 12M 61.9 Class. Human behaviour
@@
1479 hill-valley $12$ 100 122K 0.0 Class. Deterministic and simulated
@@ C.1 Architecture Details
The model architecture (see Figure 2(b)) is comprised of input embedding functions, multiple transformer encoder layers, and task-specific output heads. The key architectural parameters are summarized in Tables˜C.1 and C.2.
@@
Preprocessing   We deliberately do minimal pre-processing of the data to ensure that our approach has wide applicability. All columns containing non-numerical values are mapped to integers using scikit-learn’s [49] LabelEncoder function. The table is then standardized to 0 mean and unit variance, and outliers beyond $10$ are clipped. After retrieval, we obtain a local context $X_{\text{ctx}}$ and their labels $y_{\text{ctx}}$. $X_{\text{ctx}}$ is standardized before the forward pass to avoid distribution shifts, and $y_{\text{ctx}}$ is also standardized for the same reason if it is a regression target.
@@ Appendix E Elo and Glicko2 Ratings
In Figures˜1(a) and 1(b), we report the Elo and Glicko2 scores respectively. The results are consistent between the two plots, with TabDPT performing best on both metrics, followed by the leading TFM baseline TabPFN v2.
@@ F.1 Additional Results by Dataset Statistics
Appendix G Critical Difference Diagrams
Critical difference diagrams computed over CC18 and CTR23 fold results are given in Figures G.1, G.2, G.3, and G.4. The difference in confidence intervals compared to Table 1 are due both to how we compute the ranks and their confidence (using IQM, based on the recommendations from [1]) which computes uncertainty over different realizations of the suite. In contrast, the critical diagrams compare the uncertainty, with multiple hypothesis testing adjustments, over individual dataset realizations.
112.52.5445.55.5778.58.51010TabDPTTabMTabPFN v2MLP-PLRTabRXGBoostCatBoostLightGBMTabPFN (kNN)MLP
Figure G.1: AUC critical difference diagram
112.52.5445.55.5778.58.51010TabDPTTabPFN v2TabMTabRXGBoostMLP-PLRCatBoostLightGBMTabPFN (kNN)MLP
Figure G.2: Accuracy critical difference diagram
1122334455667788TabDPTTabPFN v2TabMTabRMLP-PLRXGBoostLightGBMCatBoost
Figure G.3: Correlation critical difference diagram
1122334455667788TabDPTTabMTabPFN v2MLP-PLRTabRXGBoostLightGBMCatBoost
Figure G.4: R2 critical difference diagram